Compute the ACT/365 interest rate for 99 days where the ACT/360 interest rate is 2.99%. Give your answer to four decimal places. Enter 4.1234% as 4.1234.
Answer #1

Given, ACT/360 for 99 days = 2.99%

For a principal of $100, Interest for 99 days= $100*2.99*99/(360*100) = $0.82225

Let the interest rate at ACT/365 be R

Then, Interest on $100 for 99 days= $100*R*99/(365*100) = $0.82225

R*99/365 =0.82225. R*0.271233= 0.82225

Therefore, Rate of interest at ACT/365 (R) = 0.82225/0.271233 = 3.031528%